# arXiv:2506.xxxxx

### v1

@@ 4.2.1 Class-I: Standard cosmological observables
$\hskip-42.67912ptp=8/3,\qquad\chi({\rm CY})=-224,\qquad\eta_{0}=6,\qquad\sigma_{0}=-4,\qquad g_{s}=0.3,$
@@ 4.2.2 Class-II: ACTivated cosmological observables
After solving the evolution equations (2.32) the canonical field evolutions for $\phi^{1}$, $\phi^{2}$ and $\phi^{3}$ are shown in figure 9, figure 10 and figure 11 respectively. As anticipated we observe see that the two fields $\phi^{2}$ and $\phi^{3}$ have identical evolution as well as minimum due to the underlying symmetry of the CY threefold itself. Moreover, using the evolution of various fields $\phi^{a}(N)$ the full scalar potential (4.13) evolves as shown in figure 12.
@@ 4.3 Demonstrating the assisted nature of inflation
which shows that the overall volume gets only a negligible shift whilst the other two fields assist during the inflation. Note that we have exploited the underlying exchange symmetry $1\leftrightarrow 2$. This clearly follows from the relation $\Delta\phi^{n}\simeq\Delta\phi/\sqrt{n}$ for $n=2$. Recall that, in the single-field and two-field version of this model, the total inflaton shift has been $\Delta\phi\simeq 5.3$.
Finally let us recall that there can be different choices of the canonical fields leading to ${\cal G}_{ab}=\delta_{ab}$ as given by $\{\varphi^{a}\}$ in (4.17) and $\{\psi^{a}\}$ in (4.18). These lead to the following shifts in the individual inflaton directions
$\Delta\varphi^{1}\simeq 4.39907,\quad\Delta\varphi^{2}\simeq 2.11929,\quad\Delta\varphi^{3}\simeq 2.11929,\qquad\Longrightarrow\,\,\Delta\varphi\simeq 5.32,$
@@
$\Delta\psi^{1}\simeq 0.0927,\quad\Delta\psi^{2}\simeq 4.61,\quad\Delta\psi^{3}\simeq 2.66,\qquad\Longrightarrow\,\,\Delta\psi\simeq 5.32,$
which clearly show that assisted two-field inflation needs smaller field excursion in the individual directions to achieve the same “effective” inflation shift $\Delta\phi=\Delta\psi=\Delta\varphi$ needed for consistently producing the cosmological observables. Note that the canonical basis $\{\varphi^{a}\}$ as defined in (4.17) does not involve the overall volume modulus and therefore all the three moduli are significantly shifted from the minimum during the inflation.
@@ 4.4 Stability of assisted fibre inflation models
This apparently shows that there is no clean hierarchy among the various individual contributions to the scalar potential, in the sense of their origin from $\alpha^{\prime}$-series or string-loop $g_{s}$ series. However we note that the hierarchy is maintained for the purpose of the so-called iterative moduli stabilization. For example, the pieces in the first line of (4.29 correspond to the stabilization of the overall volume modulus using the leading order effects in perturbative LVS. From the figure 18, we see that $V_{\rm pLVS}=V_{\alpha^{\prime}}+V_{\rm logloop}$ piece is negative during the inflationary evolution of the volume modulus and an approximately equal positive contribution from the uplifting piece compensates this to give a nearly flat de-Sitter minimum.
@@
From (4.29) we also see that that the BBHL’s $\alpha^{\prime}$-corrections which are tree-level effects in string-loop series are slightly smaller than the log-loop effects which are $g_{s}$ corrections at 1-loop level. Also, the winding type 1-loop corrections are smaller than the BBHL as well as log-loop effects. However, given the nature of the model which synthesizes several corrections from different perturbative series expansions for stabilizing the same set of moduli and building a large field inflationary model, the presence of a significant hierarchy among various individual scalar potential contributions arising from different series is anyways not likely to be anticipated/present.
However, one indeed finds that $\left|\langle V_{\rm pLVS}\rangle\right|>\left|\langle V_{g_{s}}^{W}\rangle\right|>\left|\langle V_{{\rm F}^{4}}\rangle\right|$, and subsequently the winding-type string-loop effects and the higher derivative F4 corrections are used to stabilize the remaining two moduli. These two sub-leading contributions are nearly same and opposite as nicely seen from the evolution presented in the figure 19. Further, the following eigenvalues of the mass-squared matrix indeed ensure a mass hierarchy among the various moduli
@@
where the heaviest eigenstate corresponds to the overall volume modulus ${\cal V}$ while the other ones are some combinations of all the three moduli $\{{\cal V},t^{2},t^{3}\}$. A clean mass hierarchy between the overall volume mode, e.g. see (4.27), and the two inflaton fields ensures that there is a very small shift in ${\cal V}$ modulus during inflation and the three-field evolution is effectively a two-field assisted inflationary dynamics as we have discussed.
@@ 5 Summary and Conclusions
We considered the type IIB superstring compactifications using an orientifold of a K3-fibred Calabi Yau threefold with toroidal-like volume ${\cal V}=2\,t^{1}t^{2}t^{3}$. Following the two-step moduli stabilization strategy, the complex-structure moduli and the axio-dilaton are fixed in the supersymmetric minimum by using the superpotential induced by the $F_{3}/H_{3}$ fluxes. The second step used a combination of the so-called log-loop effects [37] and the BBHL corrections [19] leading to an exponentially large volume AdS minimum [37, 48]. Subsequently, the presence of additional contributions to the scalar potential arising from the KK and winding type string-loop effects [33, 90, 32, 35] and the higher-derivative F4-corrections [30] can facilitate slow-roll inflationary dynamics. The present work can be summarized along the following steps:
@@
  • Step 1: For achieving an effective single-field inflationary model, we fixed a combination of the volume moduli, namely $t^{2}=t^{3}$ by turning on appropriate fluxes. Subsequently, the model simplifies to a two-field case, and after fixing the overall volume ${\cal V}$ in perturbative LVS, one gets a single-field fibre inflation [64]. This model have similar cosmological predictions as the standard fibre inflation [54, 55, 53, 58] which faces some strong challenge in the form of the inflaton field range bounds arising from the Kähler cone conditions [105]. However, unlike these models, the current setup does not typically face this challenge due to the absence of the rigid 4-cycle.
@@
    In this work we revisited the single-field analysis presented in [64] by numerically tracking the evolution of the inflaton field and the subsequently producing the cosmological observables. In addition, we presented additional numerical benchmark models consistent with recent ACT observations. We have found that one typically needs field excursions $\Delta\phi\simeq 6$ Mp in order to consistently produce the cosmological observables. We also observe that the analogous ACTivated models need slightly smaller field excursions.
@@
  • Step 2: There are several types of corrections arising from different series expansions which are being used for moduli stabilization and driving an effective single-field inflation. Therefore, in order to check the robustness of the assumption made in the single-field approximations we have performed a detailed numerical analysis for a two-field model, with $\{{\cal V},t^{2}\}$ moduli, and have tracked their evolution with respect to the number of e-folds. We have discussed the viability of normal and ACTivated models with consistently producing the cosmological observables.
@@
  • Step 3: From the single-field and two-field analysis we observe that one needs $\Delta\phi\simeq 6$ Mp for a successful inflationary dynamics while the overall volume remains seated at its perturbative LVS minimum. This means that even the two-field analysis in Step 2 is effectively not a multi-field inflation.
@@
For our numerical benchmark models we used multiple uplifting mechanisms with an uplifting term: $V_{\rm up}\propto{\cal V}^{-p}$ where $p=4/3$ for anti-D3 uplifting, $p=2$ for D-term uplifting and $p=8/3$ for T-brane uplifting. We showed that all these uplifting schemes can produce single and multi-field fibre inflation models with $g_{s}\simeq 0.3$, $W_{0}\simeq 5$ and ${\cal V}\simeq 10^{3}$ along with correct cosmological predictions. In typical fibre inflation models based on Swiss-Cheese Calabi Yau, the Kähler cone conditions require that ${\cal V}\gtrsim 10^{4}$ for having a sufficiently large inflaton field range [57], and subsequently one needs larger $W_{0}$ in order to match with the experimental value of the scalar perturbation amplitude. However we also note that increasing $W_{0}$ too much may enhance the gravitino mass such that it becomes heavier than the lightest bulk KK mode. Attempts for realizing fibre inflation in Swiss-Cheese based models with $W_{0}\simeq 5$ and ${\cal V}\simeq 10^{3}$ may result in pushing one or several Kähler moduli close to the boundary of the Kähler cone, which has its own new challenges regarding EFT control [105]. We also note that VEV of $W_{0}$ is intertwined with total D3 tadpole charge $Q_{D3}$ allowed by the chosen Calabi Yau orientifold [100, 64, 58]. The Kähler cone constraints are typically absent or milder in perturbative LVS model [64] and the issues with large $W_{0}\simeq{\cal O}(10)$ could be resolved by considering suitable Calabi Yau threefolds with larger $h^{1,1}$ [106, 107]. So it is very well possible that these challenges could be just because of the simple choice of Calabi Yau threefolds. However, the multi-field models will be natural and inevitable for setups with large $h^{1,1}$ as finding out an effective single-field potential in the presence of a huge variety of scalar potential contributions and maintaining the needed iterative hierarchy would be equally challenging.
@@
In our systematic analysis, we have shown that the minimal assisted fibre inflation proposal can be embedded with a range of possible uplifting schemes in perturbative LVS. We performed a detailed numerical analysis and provided several numerical benchmarks models consistent with Planck and ACT data, including the spectral tilt $n_{s}$, the running of spectral tilt $\alpha_{s}$ and the tensor-to-scalar ratio $r$. We also found that typically the tensor to scalar ratio $r$ reduces for higher values of $n_{s}$ in ACTivated models while the values of $\alpha_{s}$ can be found to be positive or negative both, with a slight change of model dependent parameter. Thus, as anticipated, multi-field inflation model is more flexible in accommodating the cosmological predictions.
@@
Finally, we discussed the viability and robustness of the assisted fibre inflation where we presented a detailed analysis of the evolution of various scalar potenial pieces during the entire inflationary period. We also discussed the issue of hierarchy among the various mass scales as in (4.31) where, we noted that the heaviest KK modes approach the string scale towards the end of inflation. This suggests that the model maybe near the edge of what is allowed in a consistent quantum gravity theory. While this does not automatically imply that the EFT breaks down [104], a careful analysis of the backreaction, field excursions, and UV completions is necessary to determine the robustness of this (or any other large field) inflationary model. Alternatively, it may require more fine tuning of our parameter space to achieve appropriate VEVs for the volume moduli such that all the KK scales remain significantly below the string mass scale without lowering (one of) them below the gravitino mass. Furthermore, it would be interesting to investigate if assisted fibre inflation could be realized in the presence of del-Pezzo divisors using K3-fibred Calabi Yau threefolds with Swiss-Cheese structure. Also it would be interesting to extend the assisted fibre inflation proposal to check the relation $\Delta\phi_{n}=\Delta\phi/\sqrt{n}$ for models with more than two inflatons. We hope to present these (and related) analysis in a future work [108].

### v2

@@ 4.2.1 Class-I: Standard cosmological observables
$\hskip-42.67912ptp=8/3,\qquad\chi({\rm CY})=-224,\qquad\eta_{0}=6,\qquad\sigma_{0}=-4,\qquad g_{s}=0.3,$ (4.24)
@@ 4.2.2 Class-II: ACTivated cosmological observables
After solving the evolution equations (2.32) the canonical field evolutions for $\phi^{1}$, $\phi^{2}$ and $\phi^{3}$ are shown in figure 9, figure 10 and figure 11 respectively. As anticipated we observe that the two fields $\phi^{2}$ and $\phi^{3}$ have identical evolution as well as minimum due to the underlying symmetry of the CY threefold itself. Moreover, using the evolution of various fields $\phi^{a}(N)$ the full scalar potential (4.13) evolves as shown in figure 12.
@@ 4.3 Demonstrating the assisted nature of inflation
which shows that the overall volume gets only a negligible shift whilst the other two fields assist during the inflation. Note that we have exploited the underlying exchange symmetry $2\leftrightarrow 3$. This clearly follows from the relation $\Delta\phi^{n}\simeq\Delta\phi/\sqrt{n}$ for $n=2$. Recall that, in the single- and two-field versions of this model, the total inflaton shift was $\Delta\phi\simeq 5.3$.
Finally, let us recall that there can be different choices of the canonical fields leading to ${\cal G}_{ab}=\delta_{ab}$ as given by $\{\varphi^{a}\}$ in (4.17) and $\{\psi^{a}\}$ in (4.18). These lead to the following shifts in the individual inflaton directions
$\Delta\varphi^{1}\simeq 4.39907,\quad\Delta\varphi^{2}\simeq 2.11929,\quad\Delta\varphi^{3}\simeq 2.11929,\qquad\Longrightarrow\,\,\Delta\varphi\simeq 5.32,$ (4.29)
@@
$\Delta\psi^{1}\simeq 0.0927,\quad\Delta\psi^{2}\simeq 4.61,\quad\Delta\psi^{3}\simeq 2.66,\qquad\Longrightarrow\,\,\Delta\psi\simeq 5.32,$ (4.30)
which clearly show that assisted two-field inflation needs smaller field excursion in the individual directions to achieve the same “effective” inflation shift $\Delta\phi=\Delta\psi=\Delta\varphi$ needed for consistently producing the cosmological observables. Note that the canonical basis $\{\varphi^{a}\}$ as defined in (4.17) does not involve the overall volume modulus and therefore the three moduli are significantly shifted from the minimum during inflation.
@@ 4.4 Stability of assisted fibre inflation models
This apparently shows that there is no clean hierarchy among the various individual contributions to the scalar potential, in the sense of their origin from $\alpha^{\prime}$-series or string-loop $g_{s}$ series. However, we note that the hierarchy is maintained for the purpose of the so-called iterative moduli stabilization. For example, the pieces in the first line of (4.31 correspond to the stabilization of the overall volume modulus using the leading-order effects in perturbative LVS. From the figure 18, we see that $V_{\rm pLVS}=V_{\alpha^{\prime}}+V_{\rm logloop}$ piece is negative during the inflationary evolution of the volume modulus, and an approximately equal positive contribution from the uplifting piece compensates for this to give a nearly flat de Sitter minimum.
@@
Equation (4.31) reveals two hierarchies: (i) BBHL’s tree-level $\alpha^{\prime}$-corrections are subdominant to the one-loop $g_{s}$ corrections (log-loop effects), and (ii) Winding-type one-loop corrections are further suppressed relative to both BBHL and log-loop terms. Crucially, however, our model synthesizes distinct perturbative expansions, i.e., $\alpha^{\prime}$, $g_{s}$, and winding ones, to stabilize moduli and achieve large-field inflation. This multi-source approach inherently disfavors large hierarchies between individual potential contributions, a consistency we observe numerically.
However, one indeed finds that $\left|\langle V_{\rm pLVS}\rangle\right|>\left|\langle V_{g_{s}}^{W}\rangle\right|>\left|\langle V_{{\rm F}^{4}}\rangle\right|$, and subsequently the winding-type string-loop effects and the higher derivative F4 corrections are used to stabilize the remaining two moduli. These two sub-leading contributions are nearly the same and opposite as nicely seen from the evolution presented in the figure 19. In addition, the following eigenvalues of the mass-squared matrix indeed ensure a mass hierarchy among the various moduli
@@
where the heaviest eigenstate corresponds to the overall volume modulus ${\cal V}$ while the other ones are some combinations of all the three moduli $\{{\cal V},t^{2},t^{3}\}$. A clean mass hierarchy between the overall volume mode, e.g. see (4.27), and the two inflaton fields ensures that there is only a very small shift in ${\cal V}$ modulus during inflation and the three-field evolution is effectively a two-field assisted inflationary dynamics as we have discussed.
@@ 5 Summary and Conclusions
We considered the type IIB superstring compactifications using an orientifold of a K3-fibred Calabi Yau threefold with toroidal-like volume ${\cal V}=2\,t^{1}t^{2}t^{3}$. Following the two-step moduli stabilization strategy, the complex-structure moduli and the axio-dilaton are fixed in the supersymmetric minimum by using the superpotential induced by the $F_{3}/H_{3}$ fluxes. The second step used a combination of the so-called log-loop effects [41] and the BBHL corrections [23] leading to an exponentially large volume AdS minimum [41, 52]. Subsequently, using an appropriate de Sitter uplifting mechanism along with a set of additional contributions to the scalar potential arising from the KK- and winding-type string-loop effects [37, 94, 36, 39] and the higher-derivative F4-corrections [34] can facilitate the slow-roll inflationary dynamics. The present work can be summarized along the following steps:
@@
  • Step 1: For achieving an effective single-field inflationary model, we fixed a combination of the volume moduli, namely $t^{2}=t^{3}$ by turning on appropriate fluxes. Subsequently, the model simplifies to a two-field case with ${\cal V}=2\,t^{1}(t^{2})^{2}$, and after fixing the overall volume ${\cal V}$ in perturbative LVS, one gets an effective single-field fibre inflation [68]. This model has similar cosmological predictions as the standard fibre inflation [58, 59, 57, 62] which faces some strong challenge in the form of the inflaton field range bounds arising from the Kähler cone conditions [109]. However, unlike these models, the current setup does not typically face this challenge due to the absence of the rigid 4-cycle as argued in [68].
@@
    In this work, we started with revisiting the single-field analysis presented in [68] by numerically tracking the evolution of the inflaton field, and subsequently producing the cosmological observables. In addition, we presented additional numerical benchmark models consistent with recent ACT observations. We found that one typically needed field excursions $\Delta\phi\simeq 6$ Mp in order to consistently produce the cosmological observables. We also observed that the analogous ACTivated models needed slightly smaller field excursions.
@@
  • Step 2: There are several types of corrections arising from different series expansions which are being used for moduli stabilization and driving an effective single-field inflation. Therefore, in order to check the robustness of the assumption made in the single-field approximations, we performed a detailed numerical analysis for a two-field model, with $\{{\cal V},t^{2}\}$ moduli, and tracked their evolution with respect to the number of e-folds. We discussed the viability of standard and ACTivated models by consistently reproducing the cosmological observables.
@@
  • Step 3: From the single-field and two-field analysis, we observed that one needed $\Delta\phi\simeq 6$ Mp for a successful inflationary dynamics, while the overall volume remained seated at its perturbative LVS minimum. This means that even the two-field analysis in Step 2 is effectively not a multi-field inflation.
@@
For our numerical benchmark models we used multiple uplifting mechanisms with an uplifting term: $V_{\rm up}\propto{\cal V}^{-p}$ where $p=4/3$ for anti-D3 uplifting, $p=2$ for D-term uplifting and $p=8/3$ for T-brane uplifting. We showed that irrespective of the uplifting schemes, one can produce the single- and multi-field fibre inflation models with $g_{s}\simeq 0.3$, $W_{0}\simeq 5$, and ${\cal V}\simeq 10^{3}$ along with correct cosmological predictions. In typical fibre inflation models based on Swiss-Cheese Calabi Yau, the Kähler cone conditions require that ${\cal V}\gtrsim 10^{4}$ to have a sufficiently large inflaton field range [61], and subsequently one needs a larger value of the parameter $W_{0}$ in order to match with the correct scalar perturbation amplitude. However, we also note that increasing $W_{0}$ too much may enhance the gravitino mass such that it becomes heavier than the lightest bulk KK mode. Attempts to realize fibre inflation in Swiss-Cheese-based models with $W_{0}\simeq 5$ and ${\cal V}\simeq 10^{3}$ may result in pushing one or several Kähler moduli close to the boundary of the Kähler cone, which has its own new challenges regarding EFT control [109]. We also note that the VEV of $W_{0}$ is intertwined with the total D3 tadpole charge $Q_{D3}$ allowed by the chosen Calabi-Yau orientifold [104, 68, 62]. The Kähler cone constraints are typically absent or milder in the perturbative LVS model [68] and the problems with large $W_{0}\simeq{\cal O}(10)$ could be resolved by considering the appropriate Calabi Yau threefold with larger $h^{1,1}$ [110, 111]. So it is very well possible that these challenges could be just because of the simple choice of Calabi Yau threefolds. However, multi-field models will be natural and inevitable for setups with large $h^{1,1}$, since finding an effective single-field potential in the presence of a huge variety of scalar potential contributions and maintaining the needed iterative hierarchy would be equally challenging.
@@
In our systematic analysis, we have shown that the minimal assisted fibre inflation proposal can be embedded with a range of possible uplifting schemes in perturbative LVS. We performed a detailed numerical analysis and provided several numerical benchmarks models consistent with Planck and ACT data, including the spectral tilt $n_{s}$, the running of spectral tilt $\alpha_{s}$ and the tensor-to-scalar ratio $r$. We also found that typically the tensor-to-scalar ratio $r$ reduces for higher values of $n_{s}$ in ACTivated models while the values of $\alpha_{s}$ can be found to be positive or negative both, with a slight change of model dependent parameter. Thus, as anticipated, multi-field inflation model is more flexible in accommodating the recent cosmological predictions.
@@
Finally, we discussed the viability and robustness of the assisted fibre inflation where we presented a detailed analysis of the evolution of various scalar potenial pieces during the entire inflationary period. We also discussed the issue of hierarchy among the various mass scales as in (4.33) where we noted that the heaviest KK modes approach the string scale towards the end of inflation. This suggests that the model maybe near the edge of what is allowed in a consistent quantum gravity theory. Although this does not automatically imply that the EFT breaks down [108], a careful analysis of the backreaction, field excursions, and UV completions is necessary to determine the robustness of this (or any other large field) inflationary model. Alternatively, it may require more fine tuning of our parameter space to achieve appropriate VEVs for the volume moduli such that all the KK scales remain significantly below the string mass scale without lowering (one of) them below the gravitino mass. Furthermore, it would be interesting to investigate if assisted fibre inflation could be realized in the presence of del-Pezzo divisors using K3-fibred Calabi Yau threefolds with Swiss-Cheese structure. Also it would be interesting to extend the assisted fibre inflation proposal to check the relation $\Delta\phi_{n}=\Delta\phi/\sqrt{n}$ for models with more than two inflatons. We hope to present these (and related) issues in a future work [112].